OpenStudy (anonymous):

What is the solution to the equation log 4 x - log 4 2 = 2

OpenStudy (campbell_st):

well you need to apply the log law for division, which is subtract the logs so you question can be written as \[\log_{4} (\frac{x}{2}) = 2\] next raise each side of the equation as a power of the base... this uses the log law \[a^{\log_{a}(b)} = b\] so \[4^{\log_{4}(\frac{x}{2})} = 4^2\] then \[\frac{x}{2} = 16\] just solve for x

OpenStudy (anonymous):

so x=8?

OpenStudy (campbell_st):

not quite... look at the equation, some number divided by 2 is equal to 16... whats the number..?

OpenStudy (anonymous):

oh so 32?

OpenStudy (campbell_st):

thats it...